# Generating Verifiable Chain of Thoughts from Exection-Traces

### v2

@@ 6. Conclusion and Future Work
Appendix: Verifiable CoT Data Samples
This appendix provides concrete examples of the verifiable Chain-of-Thought (CoT) data generated by our synthesis pipeline. The samples are organized to showcase the diversity of our dataset across two key dimensions: reasoning direction and problem difficulty. We include examples for:
  • Reasoning Direction: Forward-only, Backward-only, and Bi-directional conversational formats.
  • Problem Difficulty: Samples rated at levels 2 (Easy), 3 (Medium), and 4 (Hard) based on our quality-scoring mechanism.
Each sample is presented in a structured format that distinguishes the user prompt (the problem context) from the model’s generated response (the verifiable reasoning and final answer).
Appendix Sample: Forward Reasoning (Difficulty 2)
User Prompt (Forward) The ‘solution‘ function takes a list of events with associated probabilities, normalizes them if they do not sum to 1, and returns the cumulative distribution function (CDF) and the probability density function (PDF). Here’s the code: from functools import reduce def solution(events: list[tuple[str, float]]) -> tuple[list[float], list[float]]: # Sort events by probability sorted_events = sorted(events, key=lambda x: x[1]) # Extract probabilities probabilities = list(map(lambda x: x[1], sorted_events)) # Normalize if necessary total_prob = sum(probabilities) if total_prob != 1: probabilities = [p / total_prob for p in probabilities] # Calculate PDF pdf = probabilities # Calculate CDF cdf = [] cumulative_prob = 0 for prob in pdf: cumulative_prob += prob cdf.append(cumulative_prob) return cdf, pdf What output does the function return when given the input [(’a’, 0.5), (’b’, 0.5)]?
Assistant Response ### Understand The goal is to determine the output of the function solution when given the input events = [(’a’, 0.5), (’b’, 0.5)]‘. The functions purpose is to compute two lists: the probability density function (PDF) and the cumulative distribution function (CDF) of the input probabilities. The function sorts the events by their probabilities, extracts the probabilities, normalizes them if necessary, and then calculates the PDF and CDF. Key components: - Sorting the events by probability. - Extracting probabilities and normalizing them if they dont sum to 1. - Calculating the PDF (normalized probabilities). - Calculating the CDF by accumulating the probabilities. ### Plan To determine the output, I will: 1. Analyze the input and the initial operations (sorting and extracting probabilities). 2. Check if normalization is needed and perform it if necessary. 3. Calculate the PDF using the normalized probabilities. 4. Calculate the CDF by accumulating the PDF values. 5. Verify the final output by considering the logical flow and ensuring all steps are consistent. ### Execute 1. **Input and Sorting**: The input is events = [(’a’, 0.5), (’b’, 0.5)]‘. The sorting step sorts the events by the second element (probability), but since both probabilities are equal, the order remains ‘(’a’, 0.5), (’b’, 0.5)‘. 2. **Extracting Probabilities**: The probabilities extracted are ‘[0.5, 0.5]‘. 3. **Normalization Check**: The sum of probabilities is ‘0.5 + 0.5 = 1.0‘. Since the total is already 1, no normalization is needed, and the probabilities remain ‘[0.5, 0.5]‘. 4. **PDF Calculation**: The PDF is directly the list of probabilities, ‘[0.5, 0.5]‘. 5. **CDF Calculation**: - Initialize cumulative_prob = 0‘. - For the first probability ‘0.5‘, cumulative_prob = 0 + 0.5 = 0.5‘. Append ‘0.5‘ to cdf‘. - For the second probability ‘0.5‘, cumulative_prob = 0.5 + 0.5 = 1.0‘. Append ‘1.0‘ to cdf‘. - The resulting CDF is ‘[0.5, 1.0]‘. 6. **Final Output**: The function returns the tuple ‘(cdf, pdf)‘, which is ‘([0.5, 1.0], [0.5, 0.5])‘. ### Reflect The reasoning process involved analyzing each step of the function, ensuring that the probabilities were correctly extracted, checked for normalization, and used to compute the PDF and CDF. The calculations were straightforward, and the logic of accumulating probabilities for the CDF was correctly applied. The final output aligns with the expected behavior of the function, confirming that the reasoning was thorough and accurate. Predicted Output: ([0.5, 1.0], [0.5, 0.5])
Appendix Sample: Forward Reasoning (Difficulty 3)
User Prompt (Forward) The ‘solution‘ function approximates a solution to the Traveling Salesman Problem by using a greedy algorithm to find a route that visits all given locations exactly once, prioritizing closer locations (within 10 units) to the current location, and returns to the starting point. Here’s the code: def solution(locations): from math import sqrt def distance(p1, p2): return sqrt((p1[0] - p2[0]) ** 2 + (p1[1] - p2[1]) ** 2) if len(locations) < 2: return locations # Start from the first location start = locations[0] remaining = locations[1:] route = [start] while remaining: # Calculate distances from the current location current = route[-1] # Prioritize locations within 10 units using ternary operation prioritized = sorted( remaining, key=lambda loc: (distance(current, loc) <= 10, distance(current, loc)) ) # Select the closest location next_location = prioritized[0] route.append(next_location) remaining.remove(next_location) # Return to the starting point route.append(start) return route What does the function return when given the input [(0, 0), (1, 1), (2, 2), (3, 3)]?
Assistant Response ### Understand The goal is to determine the output of the solution function when given the input ‘[(0, 0), (1, 1), (2, 2), (3, 3)]‘. The function is designed to construct a route from a list of locations, starting from the first location, and then iteratively selecting the next location based on proximity, prioritizing those within 10 units. The function returns to the starting point after visiting all other locations. Key components include the distance function, the sorting mechanism for prioritizing locations, and the loop that builds the route. ### Plan 1. **Initialization**: Understand the initial setup of variables like start‘, remaining‘, and route‘. 2. **Loop Execution**: Analyze the loops behavior, focusing on how it selects the next location based on distance and prioritization. 3. **Return to Start**: Consider the final step where the route returns to the starting point. 4. **Verify**: Ensure the reasoning aligns with the expected behavior of the function. ### Execute 1. **Initialization**: - The function begins by checking if there are fewer than two locations, in which case it returns the input as is. This is not applicable here since there are four locations. - start is set to the first location ‘(0, 0)‘. - remaining is initialized with the rest of the locations ‘[(1, 1), (2, 2), (3, 3)]‘. - route starts with ‘[start]‘, i.e., ‘[(0, 0)]‘. 2. **Loop Execution**: - The loop continues as long as there are locations in remaining‘. - For each iteration, the current location is the last in route‘. - The remaining locations are sorted based on whether they are within 10 units and then by distance. - The closest location is selected and added to route‘, then removed from remaining‘. **Iteration Details**: - **First Iteration**: - current is ‘(0, 0)‘. - Distances to remaining locations: - ‘(1, 1)‘: $\(\sqrt${2}\) - ‘(2, 2)‘: $\(\sqrt${8}\) - ‘(3, 3)‘: $\(\sqrt${18}\) - All are within 10 units, so they are sorted by distance: ‘[(1, 1), (2, 2), (3, 3)]‘. - next_location is ‘(1, 1)‘, added to route‘, and removed from remaining‘. - **Second Iteration**: - current is ‘(1, 1)‘. - Distances to remaining locations: - ‘(2, 2)‘: $\(\sqrt${2}\) - ‘(3, 3)‘: $\(\sqrt${8}\) - Both are within 10 units, sorted by distance: ‘[(2, 2), (3, 3)]‘. - next_location is ‘(2, 2)‘, added to route‘, and removed from remaining‘. - **Third Iteration**: - current is ‘(2, 2)‘. - Distance to remaining location ‘(3, 3)‘ is $\(\sqrt${2}\). - next_location is ‘(3, 3)‘, added to route‘, and removed from remaining‘. 3. **Return to Start**: - After the loop, the route returns to the starting point ‘(0, 0)‘ by appending it to route‘. 4. **Final Route**: - The constructed route is ‘[(0, 0), (1, 1), (2, 2), (3, 3), (0, 0)]‘. ### Reflect The reasoning process aligns with the functions logic, ensuring that each step is consistent with the codes behavior. The function correctly constructs a route based on proximity and returns to the start, confirming the expected output. The approach considered the sorting mechanism and loop behavior, ensuring a thorough understanding of the functions execution. The final output is consistent with the expected behavior and the trace provided. Predicted Output: [(0, 0), (1, 1), (2, 2), (3, 3), (0, 0)]
Appendix Sample: Forward Reasoning (Difficulty 4)
User Prompt (Forward) The ‘solution‘ function identifies the pair of tuples with the smallest Euclidean distance from a sorted list of coordinate tuples using a divide-and-conquer approach, ensuring efficient performance with a time complexity of O(n log n). Here’s the code: def solution(coordinates): # Helper function to calculate Euclidean distance def euclidean_distance(point1, point2): return ((point1[0] - point2[0]) ** 2 + (point1[1] - point2[1]) ** 2) ** 0.5 # Recursive function to find the closest pair def closest_pair(coords): if len(coords) <= 3: return brute_force(coords) mid = len(coords) // 2 left_half = coords[:mid] right_half = coords[mid:] # Find closest pairs in each half left_closest = closest_pair(left_half) right_closest = closest_pair(right_half) # Determine the smaller distance min_distance = min(euclidean_distance(*left_closest), euclidean_distance(*right_closest)) closest_pair = left_closest if euclidean_distance(*left_closest) < euclidean_distance(*right_closest) else right_closest # Check the strip around the midpoint strip = [point for point in coords if abs(point[0] - coords[mid][0]) < min_distance] strip.sort(key=lambda x: x[1]) # Sort strip by y-coordinate for i in range(len(strip)): for j in range(i + 1, len(strip)): if (strip[j][1] - strip[i][1]) >= min_distance: break distance = euclidean_distance(strip[i], strip[j]) if distance < min_distance: min_distance = distance closest_pair = (strip[i], strip[j]) return closest_pair # Brute force approach for small number of points def brute_force(points): min_distance = float(’inf’) closest = None for i in range(len(points)): for j in range(i + 1, len(points)): distance = euclidean_distance(points[i], points[j]) if distance < min_distance: min_distance = distance closest = (points[i], points[j]) return closest # Sort coordinates by x-coordinate sorted_coords = sorted(coordinates, key=lambda x: x[0]) return closest_pair(sorted_coords) What output does the ‘solution‘ function return when given the input ‘[(1, 2), (3, 4)]‘?
Assistant Response ### Understand The solution function is designed to find the closest pair of points from a list of 2D coordinates using Euclidean distance. The function employs a recursive approach to solve this problem efficiently, leveraging a divide-and-conquer strategy. The key components of the function include: 1. **Euclidean Distance Calculation**: A helper function euclidean_distance computes the distance between two points. 2. **Recursive Closest Pair Function**: The closest_pair function recursively divides the list of points and finds the closest pair in each half. It also considers points near the dividing line to ensure the closest pair is found. 3. **Brute Force Method**: For small subsets of points (3 or fewer), the brute_force function is used to find the closest pair by checking all possible pairs. 4. **Sorting**: The input coordinates are sorted by their x-coordinates before processing. The functions goal is to return the pair of points with the smallest Euclidean distance. ### Plan To determine the output for the input ‘[(1, 2), (3, 4)]‘, we will: 1. **Analyze Initialization**: Understand how the input is processed initially, particularly the sorting step. 2. **Trace Recursive Logic**: Follow the recursive logic of closest_pair to see how it handles the input. 3. **Evaluate Base Case**: Since the input has only two points, the base case of the recursion will be reached quickly. 4. **Compute Distances**: Calculate the Euclidean distance between the two points to verify the result. 5. **Consider Edge Cases**: Although not directly applicable here, consider if the logic holds for minimal inputs. ### Execute 1. **Initialization**: The input ‘[(1, 2), (3, 4)]‘ is sorted by x-coordinates, resulting in sorted_coords = [(1, 2), (3, 4)]‘. 2. **Recursive Logic**: - The closest_pair function is called with sorted_coords‘. - Since the length of coords is 2, the function does not split further and directly proceeds to the brute_force function. 3. **Base Case Evaluation**: - In brute_force‘, the function iterates over all pairs of points. - For ‘[(1, 2), (3, 4)]‘, there is only one pair to consider: ‘(1, 2)‘ and ‘(3, 4)‘. 4. **Distance Calculation**: - Compute the Euclidean distance between ‘(1, 2)‘ and ‘(3, 4)‘: \[ $\text${distance} = $\sqrt${(3 - 1)^2 + (4 - 2)^2} = $\sqrt${4 + 4} = $\sqrt${8} = $\sqrt{2}$ \] - Since this is the only pair, it is the closest pair by default. 5. **Return Result**: - The function returns the pair ‘((1, 2), (3, 4))‘ as the closest pair. ### Reflect The reasoning aligns with the functions logic, as the input size is small enough to directly use the brute force method. The calculation of the Euclidean distance confirms that the pair ‘(1, 2)‘ and ‘(3, 4)‘ is indeed the closest pair. The functions design ensures that even with minimal inputs, the correct pair is returned. The approach is robust and generalizes well to larger inputs by using the divide-and-conquer strategy. Predicted Output: ((1, 2), (3, 4))
Appendix Sample: Forward Reasoning (Difficulty 4)
User Prompt (Forward) The ‘solution‘ function calculates the probability of an event occurring across a sequence of independent trials using binary partitioning and the inclusion-exclusion principle, efficiently handling up to 1,000 trials with memoization to ensure high precision. Here’s the code: def solution(probabilities): # Helper function to calculate the combined probability using inclusion-exclusion def combine_probs(p1, p2): return p1 + p2 - (p1 * p2) # Recursive approach with memoization to avoid redundant calculations memo = {} def recursive_prob(start, end): if (start, end) in memo: return memo[(start, end)] if end - start == 1: return probabilities[start] mid = (start + end) // 2 prob_left = recursive_prob(start, mid) prob_right = recursive_prob(mid, end) combined = combine_probs(prob_left, prob_right) memo[(start, end)] = combined return combined return recursive_prob(0, len(probabilities)) When I pass a list of probabilities to the ‘solution‘ function, does it return a combined probability of 1.0, considering it uses the inclusion-exclusion principle to calculate the probability?
Assistant Response ### Understand The solution function is designed to compute a combined probability from a list of individual probabilities using the inclusion-exclusion principle. The function includes a helper function combine_probs that combines two probabilities using the formula \( p1 + p2 - (p1 $\times$ p2) \). The main function uses a recursive approach with memoization to efficiently calculate the combined probability for the entire list. The recursion splits the list into halves until it reaches base cases of single probabilities, then combines them back up using the combine_probs function. The memoization helps avoid redundant calculations by storing results of subproblems. ### Plan To determine if the function returns a combined probability of 1.0, we need to: 1. Understand how the inclusion-exclusion principle is applied through the combine_probs function. 2. Analyze the recursive structure and how it processes the list of probabilities. 3. Consider edge cases, such as when the list contains probabilities that sum to 1 or when probabilities are 0 or 1. 4. Verify if the recursive combination of probabilities can logically result in 1.0. ### Execute 1. **Inclusion-Exclusion Principle**: The combine_probs function combines two probabilities \( p1 \) and \( p2 \) using the formula \( p1 + p2 - (p1 $\times$ p2) \). This formula accounts for the overlap between two events, ensuring that the combined probability does not exceed 1. 2. **Recursive Structure**: The function recursive_prob divides the list into smaller segments, computes the combined probability for each segment, and then combines these results. The base case is when the segment contains a single probability, which is returned directly. 3. **Edge Cases**: - If all probabilities are 0, the combined probability will be 0. - If all probabilities are 1, the combined probability will be 1. - If the list contains a mix of 0s and 1s, the result will depend on their arrangement and combination. 4. **Logical Combination to 1.0**: For the function to return 1.0, the combination of probabilities must account for all possible overlaps such that the final result is 1.0. This can happen if the probabilities are structured such that their combined effect covers the entire probability space. 5. **Mathematical Consideration**: The inclusion-exclusion principle ensures that as more probabilities are combined, the overlap is subtracted, preventing the result from exceeding 1.0. Given enough probabilities and their overlaps, it is possible for the combined probability to reach 1.0, especially if the probabilities are structured to cover all possible outcomes. ### Reflect The functions logic, using the inclusion-exclusion principle and recursive combination, can indeed result in a combined probability of 1.0, particularly if the input probabilities are structured to cover the entire probability space. The recursive approach with memoization ensures efficient computation, and the inclusion-exclusion principle prevents the result from exceeding 1.0. The trace confirms that the function returns 1.0 for the given input, aligning with the logical deduction. Predicted Output: True
Appendix Sample: Forward Reasoning (Difficulty 4)
User Prompt (Forward) The function schedules tasks by sorting them by deadline and using dynamic programming to determine the minimum start time for each task, then reconstructs and returns the schedule with task details including start times. Here’s the code: def solution(tasks): # Sort tasks by their deadline tasks.sort(key=lambda x: x[2]) n = len(tasks) # Initialize a DP table to store the minimum start time for each task dp = [[float(’inf’)] * (n + 1) for _ in range(n + 1)] dp[0][0] = 0 # Fill the DP table for i in range(1, n + 1): task_name, duration, deadline = tasks[i - 1] for j in range(i + 1): # If we don’t take the current task dp[i][j] = dp[i - 1][j] # If we take the current task if j > 0: start_time = max(dp[i - 1][j - 1], deadline - duration) dp[i][j] = min(dp[i][j], start_time + duration) # Reconstruct the solution result = [] time = dp[n][n] for i in range(n, 0, -1): if dp[i][n] != dp[i - 1][n]: task_name, duration, deadline = tasks[i - 1] start_time = max(dp[i - 1][n - 1], deadline - duration) result.append((task_name, duration, deadline, start_time)) time = start_time n -= 1 return result[::-1] When I provide the input ‘[("task1", 0, 0), ("task2", 2147483647, 2147483647)]‘ to the ‘solution‘ function, will it return ‘[(’task1’, 0, 0, 0), (’task2’, 2147483647, 2147483647, 0)]‘ as expected?
Assistant Response ### Understand The goal is to determine if the solution function, when given the input ‘[("task1", 0, 0), ("task2", 2147483647, 2147483647)]‘, will return ‘[(’task1’, 0, 0, 0), (’task2’, 2147483647, 2147483647, 0)]‘. The function appears to be solving a scheduling problem using dynamic programming (DP). It sorts tasks by their deadlines and calculates the earliest possible start time for each task, ensuring no task starts after its deadline. The DP table dp is used to store the minimum start time for scheduling a subset of tasks. The function then reconstructs the schedule to provide the start time for each task. ### Plan 1. **Initialization and Sorting**: Verify that the tasks are sorted by their deadlines. 2. **DP Table Setup**: Understand the initialization of the DP table and its purpose. 3. **DP Table Filling**: Analyze how the DP table is filled, focusing on the logic for deciding whether to include a task and calculating the start time. 4. **Reconstruction**: Examine how the solution is reconstructed from the DP table. 5. **Validation**: Check if the reconstructed solution matches the expected output. ### Execute 1. **Initialization and Sorting**: - The input tasks are sorted by their deadlines: ‘[("task1", 0, 0), ("task2", 2147483647, 2147483647)]‘. Since "task1" has an earlier deadline, the order remains unchanged. 2. **DP Table Setup**: - A DP table dp of size ‘(n+1) x (n+1)‘ is initialized with inf‘, except dp[0][0] = 0‘. This setup is for storing the minimum start time for scheduling j tasks out of the first i tasks. 3. **DP Table Filling**: - For each task, the table is filled considering two scenarios: including the task or not. - **Task 1 ("task1", 0, 0)**: - For j = 0‘: dp[1][0] = dp[0][0] = 0‘. - For j = 1‘: Calculate start_time = max(dp[0][0], 0 - 0) = 0‘. Update dp[1][1] = min(inf, 0 + 0) = 0‘. - **Task 2 ("task2", 2147483647, 2147483647)**: - For j = 0‘: dp[2][0] = dp[1][0] = 0‘. - For j = 1‘: Calculate start_time = max(dp[1][0], 2147483647 - 2147483647) = 0‘. Update dp[2][1] = min(0, 0 + 2147483647) = 0‘. - For j = 2‘: Calculate start_time = max(dp[1][1], 2147483647 - 2147483647) = 0‘. Update dp[2][2] = min(inf, 0 + 2147483647) = 2147483647‘. 4. **Reconstruction**: - Start from the last task and check if including it changes the DP value. - For i = 2‘ ("task2"): dp[2][2] != dp[1][2]‘, so include it. Calculate start_time = max(dp[1][1], 2147483647 - 2147483647) = 0‘. Add ‘(’task2’, 2147483647, 2147483647, 0)‘. - Update time = 0‘, n = 1‘. - For i = 1‘ ("task1"): dp[1][1] != dp[0][1]‘, so include it. Calculate start_time = max(dp[0][0], 0 - 0) = 0‘. Add ‘(’task1’, 0, 0, 0)‘. 5. **Validation**: - The reconstructed result is ‘[(’task2’, 2147483647, 2147483647, 0), (’task1’, 0, 0, 0)]‘. - Reverse it to get ‘[(’task1’, 0, 0, 0), (’task2’, 2147483647, 2147483647, 0)]‘, which matches the expected output. ### Reflect The reasoning process involved understanding the DP approach used for scheduling tasks and verifying each step of the logic. The reconstruction step was crucial in ensuring the output matched the expected result. The chosen method was robust, as it considered both the DP table filling and the reconstruction process. The final output aligns with the expected result, confirming the correctness of the functions logic. Predicted Output: [(’task1’, 0, 0, 0), (’task2’, 2147483647, 2147483647, 0)]
Appendix Sample: Backward Reasoning (Difficulty 4)
User Prompt (Backward) The function ‘solution‘ calculates and returns the mean, median, and mode of a non-empty list of integers, ensuring input validity through assertions. Here’s the code: def solution(numbers: list[int]) -> dict[str, float]: # Assertions to ensure valid input assert numbers, "Input list cannot be empty." assert all(isinstance(item, int) for item in numbers), "List must contain only integers." # Mean calculation total = sum(numbers) length = len(numbers) mean_result = total / length # Median calculation sorted_nums = sorted(numbers) half = length // 2 if length % 2 == 0: median_result = (sorted_nums[half - 1] + sorted_nums[half]) / 2 else: median_result = sorted_nums[half] # Mode calculation using a dictionary to count occurrences count_dict = {} for num in numbers: count_dict[num] = count_dict.get(num, 0) + 1 max_count = max(count_dict.values()) mode_result = min(key for key, value in count_dict.items() if value == max_count) # Return the computed statistics return {’mean’: mean_result, ’median’: median_result, ’mode’: float(mode_result)} What input would produce the output ’mean’: 2.0, ’median’: 2, ’mode’: 1 when passed to the function?
Assistant Response ### Understand The goal is to deduce the input(s) that produce the output ‘{’mean’: 2.0, median’: 2, mode’: 1}‘ when passed to the solution function. The function calculates three statistical measures: mean, median, and mode from a list of integers. The mean is the average of the numbers, the median is the middle value when sorted, and the mode is the most frequent number (with ties broken by the smallest number). The function includes assertions to ensure the input is a non-empty list of integers. ### Plan To deduce the input(s), I will reverse engineer the calculations for mean, median, and mode: 1. **Mean**: The mean is calculated as the sum of the numbers divided by the count. For a mean of 2.0, the sum of the numbers divided by their count must equal 2.0. 2. **Median**: The median is the middle value in a sorted list. For a median of 2, the middle value (or average of two middle values) must be 2. 3. **Mode**: The mode is the most frequent number. For a mode of 1, the number 1 must appear more frequently than any other number. I will explore different list configurations that satisfy these conditions, considering both odd and even lengths of the list. I will also ensure that the mode condition is met by having 1 as the most frequent number. ### Execute 1. **Mean Calculation**: - Let the list be ‘[a1, a2, ..., an]‘. - The mean condition is: \((a1 + a2 + ... + an) / n = 2.0\). - Therefore, \(a1 + a2 + ... + an = 2n\). 2. **Median Calculation**: - For an odd-length list, the median is the middle element. - For an even-length list, the median is the average of the two middle elements. - The median must be 2. 3. **Mode Calculation**: - The number 1 must appear more frequently than any other number. **Case 1: Odd-length list (n = 3)** - Mean: \(a1 + a2 + a3 = 6\). - Median: The middle element must be 2. - Mode: 1 must appear more frequently than any other number. Possible list: ‘[1, 2, 3]‘ - Sum: \(1 + 2 + 3 = 6\), Mean: \(6/3 = 2.0\). - Sorted: ‘[1, 2, 3]‘, Median: 2. - Mode: All numbers appear once, but 1 is the smallest, so mode is 1. **Case 2: Even-length list (n = 4)** - Mean: \(a1 + a2 + a3 + a4 = 8\). - Median: The average of the two middle elements must be 2. - Mode: 1 must appear more frequently than any other number. Possible list: ‘[1, 1, 2, 4]‘ - Sum: \(1 + 1 + 2 + 4 = 8\), Mean: \(8/4 = 2.0\). - Sorted: ‘[1, 1, 2, 4]‘, Median: \((1 + 2)/2 = 1.5\) (not valid). **Case 3: Odd-length list (n = 5)** - Mean: \(a1 + a2 + a3 + a4 + a5 = 10\). - Median: The middle element must be 2. - Mode: 1 must appear more frequently than any other number. Possible list: ‘[1, 1, 2, 3, 3]‘ - Sum: \(1 + 1 + 2 + 3 + 3 = 10\), Mean: \(10/5 = 2.0\). - Sorted: ‘[1, 1, 2, 3, 3]‘, Median: 2. - Mode: 1 appears twice, 3 appears twice, but 1 is the smallest, so mode is 1. ### Reflect The deduced inputs that satisfy the conditions are ‘[1, 2, 3]‘ and ‘[1, 1, 2, 3, 3]‘. Both produce the desired output. The given input ‘[1, 2, 3]‘ matches one of the deduced inputs, confirming the correctness of the deduction. The approach was exhaustive, considering both odd and even lengths, and ensuring all conditions for mean, median, and mode were met. Predicted Input: Plausible input 1: [1, 2, 3] Plausible input 2: [1, 1, 2, 3, 3]
Appendix Sample: Backward Reasoning (Difficulty 4)
User Prompt (Backward) The ‘invert_matrix‘ method in the ‘MatrixOperations‘ class calculates and returns the inverse of a given square matrix (up to 3x3), ensuring the matrix is square and non-singular with assert statements. Here’s the code: class MatrixOperations: def invert_matrix(self, matrix: list[list[float]]) -> list[list[float]]: n = len(matrix) assert all(len(row) == n for row in matrix), "Matrix must be square" det = self._recursive_determinant(matrix) assert det != 0, "Matrix is singular and cannot be inverted" cofactors = [[self._cofactor(matrix, i, j) for j in range(n)] for i in range(n)] adjugate = [[cofactors[j][i] for j in range(n)] for i in range(n)] inverse = [[adjugate[i][j] / det for j in range(n)] for i in range(n)] return inverse def _recursive_determinant(self, matrix: list[list[float]]) -> float: n = len(matrix) if n == 1: return matrix[0][0] if n == 2: return matrix[0][0] * matrix[1][1] - matrix[0][1] * matrix[1][0] det = 0 for c in range(n): det += ((-1) ** c) * matrix[0][c] * self._recursive_determinant(self._minor(matrix, 0, c)) return det def _minor(self, matrix: list[list[float]], i: int, j: int) -> list[list[float]]: return [row[:j] + row[j+1:] for row in (matrix[:i] + matrix[i+1:])] def _cofactor(self, matrix: list[list[float]], i: int, j: int) -> float: minor = self._minor(matrix, i, j) return ((-1) ** (i + j)) * self._recursive_determinant(minor) What input matrix would produce the inverse ‘[[0.6, -0.7], [-0.2, 0.4]]‘ when passed to the ‘invert_matrix‘ function?

### v3

@@ 6 Conclusion
Forward CoT Narration Prompt
This prompt generates forward-reasoning Chain-of-Thought rationales by narrating how the program’s execution transforms input to output using the execution trace, as described in Section 3.3.
forward_cot_narration_prompt You are an expert in code execution who explains things like a patient teacher. Given a Python function, its raw execution trace for the given input and output, and provided input/output values, explain how the input leads to the output. Use the raw trace as evidence to guide your reasoning steps and ensure alignment and accuracy, presenting your explanation as a clear, deductive sequence of steps-starting with the codes structure and logic, then logically deducing the outcome. If there are loops, dont walk through each iteration; instead, summarize their overall effect and connections like a human would. Mention the predicted output naturally near the end of your steps, after describing the key operations, not at the start. Code: ‘‘‘python {first_function} Raw Trace: ‘‘‘ {raw_trace} Given Input: {input_val} Given Output: {output_val} Output Format: <Steps> 1. [Natural first-person step] 2. [Next step] </Steps> {output}{output_val}{/output}
Backward CoT Narration Prompt
This prompt generates backward-reasoning Chain-of-Thought rationales by deducing the input from the output through reverse analysis of the execution trace, as described in Section 3.3.
backward_cot_narration_prompt You are an expert in code execution, reasoning like a detective unraveling a mystery in reverse. Given a Python function, its raw execution trace, and the known output, deduce the exact input that produced the given output. Begin with the output and the final line of the code, working backward through the code and execution trace to reconstruct state transitions and variable changes leading to the input. Use the execution trace as evidence to ground your reasoning and ensure accuracy. Your explanation must be a clear, step-by-step deduction, starting with the output and reversing through the codes logic to deduce the input. Do not mention or assume any input until the final step, focusing each step on the codes operations, trace evidence, and logical conditions without jumping to conclusions. Use the trace to guide your steps, but do not rely on any input value upfront to shortcut the process. For loops or repetitive structures, summarize their overall effect and connections, as a human would reason. Your final deduced input must match the input recorded in the execution trace. Code: ‘‘‘python {first_function} Raw Trace: ‘‘‘ {raw_trace} Given Output: {output_val} Output Format: <Steps> 1. [Begin with the output and final code line, analyzing their relationship and trace evidence] 2. [Continue backward, examining prior code and trace to deduce state changes] 3. [Further reverse step, connecting logic and conditions] ... N. [Conclude with the deduced input, tying together the reverse analysis] </Steps> {input_val}
Appendix Sample: Dual Agreement Clustering Pipeline
This sample demonstrates Stage B (Consensus-Based Code Selection) output, showing how Dual Agreement clusters solutions based on test consensus. The example traces one task from synthesis through filtering to Stage C selection.
Task Instruction: Write a Python function that computes the Greatest Common Divisor (GCD) of two positive integers using the Euclidean algorithm. The function should efficiently handle large numbers and return the GCD as an integer.
Stage A Output (Input to Stage B): For this task instruction, Stage A synthesis produces:
  • Candidate Solutions: 5 independently generated code implementations
  • Candidate Tests: 25 independently generated test cases
For this task instruction, Stage A synthesis produces 5 candidate solutions and 25 candidate tests.
Stage B Processing: Execute All $5\times 25=125$ Solution-Test Pairs
After executing all pairs and grouping solutions by identical test outcomes, we obtain multiple clusters. We show two representative clusters that emerge from this process.
Cluster 1 (High Consensus - SELECTED for Stage C)
Solutions in this cluster: 4 out of 5 generated implementations Tests passing for all 4 solutions: 18 out of 25 generated tests
Representative Solutions from Cluster 1 Solution 1 (Iterative - Assignment-based): def solution(a, b): while b: a, b = b, a % b return a Solution 2 (Iterative - Temporary variable): def solution(a, b): while b != 0: temp = b b = a % b a = temp return a Solution 3 (Recursive): def solution(a, b): if b == 0: return a return solution(b, a % b) Solution 4 (Modulo with abs): def solution(a, b): a, b = abs(a), abs(b) while b: a, b = b, a % b return a Tests All 4 Solutions Pass (18 out of 25, representative shown): def test_basic_case(): assert solution(48, 18) == 6 def test_large_numbers(): assert solution(1071, 462) == 21 def test_coprime_numbers(): assert solution(17, 19) == 1 def test_multiple_common_factors(): assert solution(360, 240) == 120 # ... 14 more tests covering edge cases, # large values, and reversed arguments
Cluster 1 Metrics:
  • Solution Agreement: 4 solutions with identical test pass patterns
  • Test Consensus: 18 tests consistently passed by all 4 solutions